OpenStudy (anonymous):

What is the probability of not getting a total of 3 when two die are rolled?

OpenStudy (anonymous):

5/6

OpenStudy (unklerhaukus):

there are thirty six possible rolls only two of which sum to 3 the probabilty of not getting a sum of three is\[1-P(3)=1-\frac{2}{36}\]

OpenStudy (ajprincess):

 the circled one is the probability of getting a sum of 3. probability of nt getting sum of three is =1-2/36